4479:

A certain radioactive atom has 90 protons and 142 neutrons. If this atom undergoes alpha decay, what is the mass number of the daughter atom? The mass number of the daughter atom is ___________.

ImLookingToCheat:

The mass number of the daughter atom is 228

4479:

@imlookingtocheat wrote:
The mass number of the daughter atom is 228
Are you completely 100% sure?

ImLookingToCheat:

Yes im 100% sure

MusicGeek:

I think it's 232.

4479:

@musicgeek wrote:
I think it's 232.
show evidence please

MusicGeek:

wait. nvm it is 228.

MusicGeek:

cuz 232 = x + 4 which goes to x = 232 - 4 = 228 which makes 228 the answer.

4479:

@musicgeek wrote:
cuz 232 = x + 4 which goes to x = 232 - 4 = 228 which makes 228 the answer.
thx <3

MusicGeek:

@4479 wrote:
@musicgeek wrote:
cuz 232 = x + 4 which goes to x = 232 - 4 = 228 which makes 228 the answer.
thx <3
yw